If an 0.860 m aqueous solution freezes at –2.10 °C, what is the van\'t Hoff factor, i, of the solute? Kf values can be found here.
Answer #1

depression in freezing ponit =kf*m*i

kf= Freezing point depression constant= 1.86 0degc/m

m= molality =0.86

2.1 = 1.86*0.86*i where i = Van't Hoff factor

i= 2.1/(1.86*0.86)=1.3. This can be approximated to 1